OpenStudy (marcelie):

help please Find the slope of the tangent line to the given curve at the point corresponding to the specified value of the parameter

OpenStudy (marcelie):

\[x=\ln t\] \[y=1+t^2\] t= 1

OpenStudy (marcelie):

x = ln t y= 1+t^2 x= ln (1) y=1+ (1)^2 x= 0 y= 2 (0,2)

OpenStudy (marcelie):

@ganeshie8

OpenStudy (marcelie):

\[\frac{ dy }{ dx } 1+t^2\] ____ = dy/dx = 2t^2 \[\frac{ dx }{ dt } \frac{ 1 }{ t }\]

OpenStudy (marcelie):

not sure if i did it right..

OpenStudy (caozeyuan):

\[\frac{ dy }{ dx }=\frac{ dy }{ dt }*\frac{ dt }{ dx }\]

OpenStudy (marcelie):

oh yea

OpenStudy (caozeyuan):

\[\frac{ dt }{ dx }=e ^{x}\]

OpenStudy (caozeyuan):

\[\frac{ dy }{ dt }=2t\]

OpenStudy (caozeyuan):

t=1, so dy/dt=2 and dt/dx=t=1

OpenStudy (caozeyuan):

so answer is 2

OpenStudy (irishboy123):

\(\dfrac{dy}{dx} = \dfrac{\dot y}{\dot x} = \dfrac{2t}{\frac{1}{t}} = 2t^2\)

OpenStudy (irishboy123):

plug in the t value

OpenStudy (reemii):

You can also reason with the geometrical interpretation (pretty much the same as what the others showed above): The velocity of the point following the curve is: \((x'(t), y'(t))\) (velocity is a vector).  So: \((x'(t),y'(t)) = (\frac1t, 2t)\). The slope is: \(\frac{\Delta y}{\Delta x} = \frac{2t}{1/t}= 2t^2\). Pretty much the same calculations.

OpenStudy (reemii):